OpenStudy (anonymous):

What is the factored form of this? d^2 + 12d + 32

OpenStudy (anonymous):

a + b = 12 a * b = 32

OpenStudy (anonymous):

How do I factor it? I don't understand.

OpenStudy (anonymous):

8 + 4 = 12 8 * 4 = 32

OpenStudy (anonymous):

(d+8)(d+4)

OpenStudy (anonymous):

(d+ a)(d+ b)

OpenStudy (anonymous):

d^2 + 12d + 32. Factoring is likely solving the quadratic equation d^2 + 12d + 32 = 0. When a = 1, the fastest way is using the Diagonal Sum Method. Both roots are negative. Compose factor pairs of c = 32. Proceeding: (-1, -32)(-2, -16)(-4, -8). This last sum is : -4 - 8 = -12 = -b. Then, the 2 real roots are: -4, and -8. Factoring form: (d + 4)(d + 8).

OpenStudy (anonymous):

Would d^2 - 12d + 32 = (d - 4) (d + 8) ?

OpenStudy (anonymous):

@thu1935

OpenStudy (anonymous):

Factor d^2 - 12d + 32 (d - 4) (d + 8) right?

OpenStudy (anonymous):

yes